OpenStudy (anonymous):

**MEDAL AWARDED FOR BEST RESPONSE*** Mia wants to deposit $5,535 into savings accounts at three different banks: Southern Bank, 3rd State Bank, and Clutch Bank. She will deposit four times as much into Southern Bank as 3rd State Bank because they offer a higher interest rate. She also expects the Clutch Bank deposit to be 35% of the total of the other two deposits. How much will Mia deposit into the Clutch Bank savings account? $3,280 $820 $1,640 $1,435 Also, please don't copy and paste an answer from yahoo, thanks :)

OpenStudy (anonymous):

let x = amount deposited into the Clutch Bank y = 3rd State Bank deposit 4y = Southern bank deposit x = 5535 - 5y and x = 0.35(5y) 5535 - 5y = 0.35(5y) 5535 - 5y = 1.75y 5535 = 6.75y y = 820 x = 1435

OpenStudy (anonymous):

Yahoo?

OpenStudy (anonymous):

yep !

OpenStudy (anonymous):

lol I didn't read Also, please don't copy and paste an answer from yahoo, thanks :)

OpenStudy (anonymous):

Hahah okay, thanks anyways!